# Dynamic Novel View Synthesis in High Dynamic Range

### v2

@@ Abstract
High Dynamic Range Novel View Synthesis (HDR NVS) seeks to learn an HDR 3D model from Low Dynamic Range (LDR) training images captured under conventional imaging conditions. Current methods primarily focus on static scenes, implicitly assuming all scene elements remain stationary and non-living. However, real-world scenarios frequently feature dynamic elements, such as moving objects, varying lighting conditions, and other temporal events, thereby presenting a significantly more challenging scenario. To address this gap, we propose a more realistic problem named HDR Dynamic Novel View Synthesis (HDR DNVS), where the additional dimension “Dynamic” emphasizes the necessity of jointly modeling temporal radiance variations alongside sophisticated 3D translation between LDR and HDR. To tackle this complex, intertwined challenge, we introduce HDR-4DGS, a Gaussian Splatting-based architecture featured with an innovative dynamic tone-mapping module that explicitly connects HDR and LDR domains, maintaining temporal radiance coherence by dynamically adapting tone-mapping functions according to the evolving radiance distributions across the temporal dimension. As a result, HDR-4DGS achieves both temporal radiance consistency and spatially accurate color translation, enabling photorealistic HDR renderings from arbitrary viewpoints and time instances. Extensive experiments demonstrate that HDR-4DGS surpasses existing state-of-the-art methods in both quantitative performance and visual fidelity. Source code will be released.
@@ 1 Introduction
Recent years have seen remarkable progress in Novel View Synthesis (NVS), which reconstructs 3D scene representations from multi-view images to enable photorealistic rendering from arbitrary viewpoints. These advances have fueled applications in gaming (Niedermayr et al., 2024), AR / VR (Zhou et al., 2018), and autonomous driving (Wang et al., 2024a). However, most existing NVS methods (Gao et al., 2022; Wu et al., 2024c) operate under two limiting assumptions: static scenes and low dynamic range inputs. These constraints significantly hinder their applicability in real-world environments, which often exhibit complex motion, time-varying illumination, and sensor-imposed luminance clipping.
@@
Dynamic Novel View Synthesis (DNVS) extends NVS to scenes with temporal dynamics, reconstructing 4D radiance fields coherent in both space and time. Recent efforts in DNVS (Fang et al., 2022; Gao et al., 2021; Tian et al., 2023; Duan et al., 2024; Yang et al., 2024b; Cho et al., 2024) have made notable progress in modeling dynamic geometry and appearance. However, they remain restricted to LDR imagery, which fails to capture the full spectrum of scene radiance. As a result, these methods struggle under high-contrast conditions (e.g., direct sunlight or low-light environments) where overexposure and underexposure lead to significant information loss. More fundamentally, LDR imaging is inherently misaligned with human visual perception, even under moderate lighting conditions. While the human eye can adapt to luminance levels spanning over ten orders of magnitude (Reinhard, 2020), typical LDR sensors cover only a narrow dynamic range. In addition, LDR images compress radiance through nonlinear camera response functions (CRFs), distorting local contrast, suppressing brightness gradients, and reducing color fidelity. These limitations not only impair perceptual quality under extreme lighting but also degrade the realism of ordinary scenes.
@@
High Dynamic Range (HDR) imaging addresses these shortcomings by capturing a significantly broader range of luminance and color. By preserving detail in both highlights and shadows and maintaining fine-grained contrast, HDR techniques provide a closer match to the perceptual capabilities of the human visual system. Recent works in HDR NVS (Huang et al., 2022; Cai et al., 2024; Liu et al., 2025) have attempted to reconstruct HDR content from multi-exposure LDR images of calibrated sensors. However, these methods are restricted to static scenes, implicitly assuming that all elements remain fixed over time. In contrast, real-world HDR scenarios are inherently dynamic, often involving moving objects, shifting illumination, and transient phenomena. These dynamics violate the assumptions of existing HDR NVS pipelines and introduce significant challenges: non-rigid motion and temporal variation create complex spatiotemporal inconsistencies, while the lack of reliable luminance priors from sparse LDR observations leads to severe photometric ambiguities.
@@
Our contributions are: (I) We formally introduce the HDR DNVS problem for the first time, which requires learning 4D HDR radiance fields with dynamic geometry and temporally coherent appearance from sparse LDR video input. (II) We propose HDR-4DGS, a novel Gaussian Splatting architecture featuring dynamic tone-mapping for adaptively bridging the LDR and HDR domains under complex spatiotemporal variations. (III) To enable rigorous quantitative evaluation of HDR DNVS methods, we introduce HDR-4D-Syn and HDR-4D-Real — two novel benchmark datasets comprising 8 high-fidelity synthetic scenes and 4 real-world captured sequences, respectively. Each scene is meticulously annotated with ground-truth HDR images, time-varying 3D geometry, and synchronized multi-view LDR observations. (IV) Extensive experiments show that HDR-4DGS achieves state-of-the-art performance in both quantitative metrics and perceptual quality on challenging dynamic scenes.
@@ 2 Related Work
Novel view synthesis (NVS) has seen transformative progress via neural rendering techniques. Early multi-view geometric methods (Schonberger & Frahm, 2016; Wang et al., 2021) face limitations in handling occlusions, textureless regions, and computing efficiency (Jiang, 2023). Modern approaches leverage continuous scene representations via deep networks, such as Neural Radiance Fields (NeRF) (Mildenhall et al., 2021) and the variants (Deng et al., 2022; Roessle et al., 2022; Wei et al., 2021; Xu et al., 2022) establishing coordinate-space implicit modeling for photorealistic synthesis under spatial smoothness constraints. However, NeRF’s ray-marching paradigm suffers from high computational costs (Luo et al., 2024). In contrast, 3DGS (Kerbl et al., 2023) introduces an efficient point-based representation by parameterizing scenes as anisotropic 3D Gaussians, decoupling geometry and appearance while enabling real-time rendering with complex visual effects (e.g., specular highlights). Recent advances enhance 3DGS through frequency-domain supervision (Liang et al., 2024), depth-regularized optimization (Chung et al., 2024; Kung et al., 2024; Li et al., 2024), and memory-efficient designs (Wang et al., 2024b; Chen et al., 2024; Fan et al., 2024; Lu et al., 2024; Yang et al., 2025), achieving superior rendering quality. Nevertheless, these methods predominantly focus on static LDR scenes, limiting their applicability to real-world dynamic scenarios.
HDR novel view synthesis (HDR NVS) aims to reconstruct HDR scenes from multi-view LDR images. Early works like HDR-NeRF (Huang et al., 2022) extend neural radiance fields by incorporating an MLP-based tone-mapping module to bridge physical radiance and digital color spaces. To realize real-time rendering, HDR-GS (Cai et al., 2024) introduces a 3DGS framework with a neural tone-mapper that explicitly models HDR-to-LDR radiance transformations, achieving real-time HDR rendering while surpassing NeRF-based quality. Recently, GaussHDR (Liu et al., 2025) proposes to unify 3D and 2D tone mapping in 3D Gaussian Splatting to facilitate HDR rendering. However, a critical limitation of these methods is their exclusive focus on static scenes; none explicitly model temporal dynamics, both in spatial and color space. Consequently, they are fundamentally unable to address the core challenges posed by real-world scenarios involving time-varying geometry, non-rigid motion, or temporally evolving illumination.
@@
Dynamic novel view synthesis (DNVS) focuses on modeling dynamic scenes with time-varying geometry and radiance. The key challenge lies in capturing intrinsic spatiotemporal correlations. Building upon NeRF (Mildenhall et al., 2021), two primary approaches have emerged: i) implicit / explicit spatiotemporal representations decompose scenes into time-aware feature grids to learn 6D plenoptic functions (Cao & Johnson, 2023; Li et al., 2022; Fridovich-Keil et al., 2023; Wang et al., 2023), and ii) deformation-aware fields model motion through deformable transformations (Pumarola et al., 2021; Song et al., 2023; Abou-Chakra et al., 2024). Recent advances leverage 3DGS for dynamic rendering along two directions: i) deformation-based models (Wu et al., 2024b; Kratimenos et al., 2024; Bae et al., 2024; Shan et al., 2025) maintain canonical Gaussians deformed via time-varying fields, trading precise motion tracking for temporal continuity; ii) hyper-dimensional representations (Yang et al., 2024b; Duan et al., 2024) extend Gaussians to 4D by introducing temporal centers and spatiotemporal rotations. However, all these methods are limited to LDR outputs, and we extend this to high-fidelity HDR DNVS problem.
@@ 3 Method
Problem In HDR DNVS, we aim to learn an HDR 3D model $\mathcal{F}$ for a target dynamic scene $G$, $\mathcal{F}:(t^{\prime},V^{\prime})\rightarrow\mathbf{I}_{t^{\prime},V^{\prime}}^{h}$, that would render an HDR image $\mathbf{I}_{t^{\prime},V^{\prime}}^{h}$ for any timestamp $t^{\prime}$ and viewpoint $V^{\prime}$. To that end, we capture a set of multi-exposure LDR training images $\mathbf{I}^{l}=\{\mathbf{I}_{1}^{l},\cdots,\mathbf{I}_{t}^{l},\cdots,\mathbf{I}_{T}^{l}\}$, each $\mathbf{I}_{t}^{l}$ associated with the exposure time $e_{t}$ sampled from the choices $E=\{e_{1},\cdots,e_{P}\}$, and the camera viewpoint $V_{t}$ selected from $Q$ distinct viewpoints $\mathbf{V}=\{V_{1},\cdots,V_{Q}\}$ where $T$ denotes the total timesteps and $P$ the number of exposure time choices. We may have optional access to coupled HDR training data $\mathbf{I}^{h}=\{\mathbf{I}_{1}^{h},\cdots,\mathbf{I}_{T}^{h}\}$.
@@ 3.2 Dynamic scene representation
HDR-4DGS can integrate with existing dynamic scene representation to better capture spatiotemporal variations. Specifically, we adopt the 4D Gaussian Splatting (4DGS) framework (Yang et al., 2024a) due to its conceptual elegance and coherent formulation of dynamic scenes.
@@ 3.3 Dynamic Tone Mapper
where $d$ denotes the dimension of this context embedding. DRCL can be generally realized by any existing sequence model such as RNN (Salehinejad et al., 2017), LSTM (Staudemeyer & Morris, 2019), GRU (Cho et al., 2014) or Transformer (Vaswani et al., 2017). We show that this design choice is not sensitive with only marginal effect to the overall performance (see Tab. 4.3).
@@ 3.4 Model optimization
To mitigate overfitting associated with directly applying 3D tone mapping on HDR Gaussian fields (Liu et al., 2025), we adopt an extra pixel-level supervision in addition to existing ray-level supervision over the dynamic tone mapper, which introduces additional constraints to facilitate robust CRF learning: $\mathbf{I}_{t,2d}^{l}=g_{\theta}([\log\ \mathbf{I}_{t,2d}^{h}+\log\ e_{t},\mathbf{f}_{t}])$, where $\mathbf{I}_{t,2d}^{h}$ is the HDR image rasterized by the HDR Gaussian model, and $\mathbf{I}_{t,2d}^{l}$ is the tone-mapped LDR image at time $t$. During training, we supervise both the tone-mapped LDR image $\mathbf{I}_{t,2d}^{l}$ and the LDR image $\mathbf{I}_{t,3d}^{l}$ rendered directly from the LDR model. This dual supervision improves the learned tone mapper’s generalization capability, as validated in our ablation (Tab. 4.3).
We define the image reconstruction loss as a weighted combination of $L_{1}$ and D-SSIM loss (Huang et al., 2022; Nilsson & Akenine-Möller, 2020):
@@
where $\mathbf{I}_{t}^{l}$ is the ground-truth LDR image. The tone-mapped HDR images $\hat{\mathbf{I}}_{t,2d}^{h}$ and $\hat{\mathbf{I}}_{t}^{h}$ are derived from the rendered HDR image $\mathbf{I}_{t,2d}^{h}$ and ground-truth HDR image $\mathbf{I}_{t}^{h}$, respectively, using $\mu$-law compression: $\mathbf{\hat{I}}^{h}=\frac{\log(1+\mu\cdot\text{norm}(\mathbf{I}^{h}))}{\log(1+\mu)}$ where $\mu$ is a compression factor and norm()($\cdot$) is min-max normalization. This transformation aligns HDR and LDR domains for consistent comparison. Notably, we choose the LDR images $\mathbf{I}_{t,3d}^{l}$ rasterized by Gaussian Splatting as our final LDR rendering results.
@@ 4 Experiments
Implementation details. HDR-4DGS is trained with the Adam optimizer using the same parameters as 4DGS (Yang et al., 2024b). For our dynamic tone mapper, the learning rate is set to $5\times 10^{-4}$, and the dimension of temporal radiance context features is set to 2. We adopt the same structure of per-channel tone-mapping functions in our dynamic tone mapper as HDR-GS (Cai et al., 2024) and GRU (Cho et al., 2014) is adopted to implement the dynamic radiance context learner by default. For equation 7, $\lambda$ is set to 0.2, and $\alpha$ is set to 0.6 if HDR ground truth is available.
Evaluation metrics. We adopt the PSNR and SSIM as quantitative metrics, LPIPS as an additional perceptual metric. Following prior works (Cai et al., 2024; Huang et al., 2022; Liu et al., 2025), we apply Photomatix Pro (HDRsoft Team, 2025) to convert HDR images into displayable LDR images for qualitative visualization and fair comparison. Futhermore, training time and inference speed (fps) are reported. Results are averaged over all scenes.

### v3

@@ Abstract
High Dynamic Range Novel View Synthesis (HDR NVS) seeks to learn an HDR 3D model from Low Dynamic Range (LDR) training images captured under conventional imaging conditions. Current methods primarily focus on static scenes, implicitly assuming all scene elements remain stationary and non-living. However, real-world scenarios frequently feature dynamic elements, such as moving objects, varying lighting conditions, and other temporal events, thereby presenting a significantly more challenging scenario. To address this gap, we propose a more realistic problem named HDR Dynamic Novel View Synthesis (HDR DNVS), where the additional dimension “Dynamic” emphasizes the necessity of jointly modeling temporal radiance variations alongside sophisticated 3D translation between LDR and HDR. To tackle this complex, intertwined challenge, we introduce HDR-4DGS, a Gaussian Splatting-based architecture featured with an innovative dynamic tone-mapping module that explicitly connects HDR and LDR domains, maintaining temporal radiance coherence by dynamically adapting tone-mapping functions according to the evolving radiance distributions across the temporal dimension. As a result, HDR-4DGS achieves both temporal radiance consistency and spatially accurate color translation, enabling photorealistic HDR renderings from arbitrary viewpoints and time instances. Extensive experiments demonstrate that HDR-4DGS surpasses existing state-of-the-art methods in both quantitative performance and visual fidelity. Source code is available at https://github.com/prinasi/HDR-4DGS.
@@ 1 Introduction
Recent years have seen remarkable progress in Novel View Synthesis (NVS), which reconstructs 3D scene representations from multi-view images to enable photorealistic rendering from arbitrary viewpoints. These advances have fueled applications in gaming (Niedermayr et al., 2024), AR / VR (Zhou et al., 2018), and autonomous driving (Wang et al., 2024b). However, most existing NVS methods (Gao et al., 2022; Wu et al., 2024c) operate under two limiting assumptions: static scenes and low dynamic range inputs. These constraints significantly hinder their applicability in real-world environments, which often exhibit complex motion, time-varying illumination, and sensor-imposed luminance clipping.
@@
Dynamic Novel View Synthesis (DNVS) extends NVS to scenes with temporal dynamics, reconstructing 4D radiance fields coherent in both space and time. Recent efforts in DNVS (Fang et al., 2022; Gao et al., 2021; Tian et al., 2023; Duan et al., 2024; Yang et al., 2024; Cho et al., 2024) have made notable progress in modeling dynamic geometry and appearance. However, they remain restricted to LDR imagery, which fails to capture the full spectrum of scene radiance. As a result, these methods struggle under high-contrast conditions (e.g., direct sunlight or low-light environments) where overexposure and underexposure lead to significant information loss. More fundamentally, LDR imaging is inherently misaligned with human visual perception, even under moderate lighting conditions. While the human eye can adapt to luminance levels spanning over ten orders of magnitude (Reinhard, 2020), typical LDR sensors cover only a narrow dynamic range. In addition, LDR images compress radiance through nonlinear camera response functions (CRFs), distorting local contrast, suppressing brightness gradients, and reducing color fidelity. These limitations not only impair perceptual quality under extreme lighting but also degrade the realism of ordinary scenes.
@@
High Dynamic Range (HDR) imaging addresses these shortcomings by capturing a significantly broader range of luminance and color. By preserving detail in both highlights and shadows and maintaining fine-grained contrast, HDR techniques provide a closer match to the perceptual capabilities of the human visual system. Recent works in HDR NVS (Huang et al., 2022; Cai et al., 2024; Liu et al., 2025a) have attempted to reconstruct HDR content from multi-exposure LDR images of calibrated sensors. However, these methods are restricted to static scenes, implicitly assuming that all elements remain fixed over time. In contrast, real-world HDR scenarios are inherently dynamic, often involving moving objects, shifting illumination, and transient phenomena. These dynamics violate the assumptions of existing HDR NVS pipelines and introduce significant challenges: non-rigid motion and temporal variation create complex spatiotemporal inconsistencies, while the lack of reliable luminance priors from sparse LDR observations leads to severe photometric ambiguities.
@@
Our contributions are: (I) We introduce the HDR DNVS problem for the first time, which requires learning 4D HDR radiance fields with dynamic geometry and temporally coherent appearance from sparse LDR input. (II) We propose HDR-4DGS, a novel Gaussian Splatting architecture featuring dynamic tone-mapping for adaptively bridging the LDR and HDR domains under complex spatiotemporal variations. (III) To enable quantitative evaluation of HDR DNVS methods, we introduce HDR-4D-Syn and HDR-4D-Real — two novel benchmark datasets comprising 8 high-fidelity synthetic scenes and 4 real-world captured sequences, respectively. Each scene is meticulously annotated with ground-truth HDR images, time-varying 3D geometry, and synchronized multi-view LDR observations. (IV) Extensive experiments show that HDR-4DGS achieves state-of-the-art performance in both quantitative metrics and perceptual quality on challenging dynamic scenes.
@@ 2 Related Work
Novel view synthesis (NVS) has seen transformative progress via neural rendering techniques. Early multi-view geometric methods (Schonberger and Frahm, 2016; Wang et al., 2021) face limitations in handling occlusions, textureless regions, and computing efficiency (Jiang, 2023). Modern approaches leverage continuous scene representations via deep networks, such as Neural Radiance Fields (NeRF) (Mildenhall et al., 2021) and the variants (Deng et al., 2022; Roessle et al., 2022; Wei et al., 2021; Xu et al., 2022) establishing coordinate-space implicit modeling for photorealistic synthesis under spatial smoothness constraints. However, NeRF’s ray-marching paradigm suffers from high computational costs (Luo et al., 2024). In contrast, 3DGS (Kerbl et al., 2023) introduces an efficient point-based representation by parameterizing scenes as anisotropic 3D Gaussians, decoupling geometry and appearance while enabling real-time rendering with complex visual effects (e.g., specular highlights). Recent advances enhance 3DGS through frequency-domain supervision (Liang et al., 2024), depth-regularized optimization (Chung et al., 2024; Kung et al., 2024; Li et al., 2024), and memory-efficient designs (Wang et al., 2024c; Chen et al., 2024; Fan et al., 2024; Lu et al., 2024; Yang et al., 2025), achieving superior rendering quality. Nevertheless, these methods predominantly focus on static LDR scenes, limiting their applicability to real-world dynamic scenarios.
HDR novel view synthesis (HDR NVS) aims to reconstruct HDR scenes from multi-view LDR images. Early works like HDR-NeRF (Huang et al., 2022) extend neural radiance fields by incorporating an MLP-based tone-mapping module to bridge physical radiance and digital color spaces. To realize real-time rendering, HDR-GS (Cai et al., 2024) introduces a 3DGS framework with a neural tone-mapper that explicitly models HDR-to-LDR radiance transformations, achieving real-time HDR rendering while surpassing NeRF-based quality. Recently, GaussHDR (Liu et al., 2025a) proposes to unify 3D and 2D tone mapping in 3D Gaussian Splatting to facilitate HDR rendering. However, a critical limitation of these methods is their exclusive focus on static scenes; none explicitly model temporal dynamics, both in spatial and color space. Consequently, they are fundamentally unable to address the core challenges posed by real-world scenarios involving time-varying geometry, non-rigid motion, or temporally evolving illumination.
@@
Dynamic novel view synthesis (DNVS) focuses on modeling dynamic scenes with time-varying geometry and radiance. The key challenge lies in capturing intrinsic spatiotemporal correlations. Building upon NeRF (Mildenhall et al., 2021), two primary approaches have emerged: i) implicit / explicit spatiotemporal representations decompose scenes into time-aware feature grids to learn 6D plenoptic functions (Cao and Johnson, 2023; Li et al., 2022b; Fridovich-Keil et al., 2023; Wang et al., 2023), and ii) deformation-aware fields model motion through deformable transformations (Pumarola et al., 2021; Song et al., 2023; Abou-Chakra et al., 2024). Recent advances leverage 3DGS for dynamic rendering along two directions: i) deformation-based models (Wu et al., 2024b; Kratimenos et al., 2024; Bae et al., 2024; Shan et al., 2025) maintain canonical Gaussians deformed via time-varying fields, trading precise motion tracking for temporal continuity; ii) hyper-dimensional representations (Yang et al., 2024; Duan et al., 2024) extend Gaussians to 4D by introducing temporal centers and spatiotemporal rotations. However, all these methods are limited to LDR outputs, and we extend this to high-fidelity HDR DNVS problem.
@@ 3 Method
Problem In HDR DNVS, we aim to learn a HDR 3D model $\mathcal{F}_{h}$ for a target dynamic scene $G$, ${\color[rgb]{0,0,0}\mathcal{F}_{h}}:(t^{\prime},V^{\prime})\rightarrow\mathbf{I}_{t^{\prime},V^{\prime}}^{h}$, that would render an HDR image $\mathbf{I}_{t^{\prime},V^{\prime}}^{h}$ for any timestamp $t^{\prime}$ and viewpoint $V^{\prime}$. To that end, we capture a set of multi-exposure LDR training images $\mathbf{I}^{l}=\{\mathbf{I}_{1}^{l},\cdots,\mathbf{I}_{t}^{l},\cdots,\mathbf{I}_{T}^{l}\}$, each $\mathbf{I}_{t}^{l}$ associated with the exposure time $e_{t}$ sampled from the choices $E=\{e_{1},\cdots,e_{P}\}$, and the camera viewpoint $V_{t}$ selected from $Q$ distinct viewpoints $\mathbf{V}=\{V_{1},\cdots,V_{Q}\}$ where $T$ denotes the total timesteps and $P$ the number of exposure time choices. We may have optional access to coupled HDR training data $\mathbf{I}^{h}=\{\mathbf{I}_{1}^{h},\cdots,\mathbf{I}_{T}^{h}\}$.
@@ 3.2 Dynamic scene representation
HDR-4DGS can integrate with existing dynamic scene representation to better capture spatiotemporal variations. Specifically, we adopt the 4D Gaussian Splatting (4DGS) framework (Yang et al., 2024) due to its conceptual elegance and coherent formulation of dynamic scenes.
@@ 3.3 Dynamic Tone Mapper
where $d$ denotes the dimension of this context embedding. DRCL can be generally realized by any existing sequence model such as RNN (Salehinejad et al., 2017), LSTM (Staudemeyer and Morris, 2019), GRU (Cho et al., 2014) or Transformer (Vaswani et al., 2017).
@@ 3.4 Model optimization
To mitigate overfitting associated with directly applying 3D tone mapping on HDR Gaussian fields (Liu et al., 2025a), we adopt an extra pixel-level supervision in addition to existing ray-level supervision over the dynamic tone mapper, which introduces additional constraints to facilitate robust CRF learning: $\mathbf{I}_{t,2}^{l}=g_{\theta}([\log\ \mathbf{I}_{t,2D}^{h}+\log\ e_{t},\mathbf{f}_{t}])$, where $\mathbf{I}_{t,2D}^{h}$ is the HDR image rasterized by the HDR Gaussian model, and $\mathbf{I}_{t,2D}^{l}$ is the tone-mapped LDR image at time $t$. During training, we supervise both the tone-mapped LDR image $\mathbf{I}_{t,2D}^{l}$ and the LDR image $\mathbf{I}_{t,3D}^{l}$ rendered directly from the LDR model. This dual supervision improves the learned tone mapper’s generalization capability, as validated in our ablation (Tab. 4.4).
We define the image reconstruction loss as a weighted combination of $L_{1}$ and D-SSIM loss (Huang et al., 2022; Nilsson and Akenine-Möller, 2020):
@@
where $\mathbf{I}_{t}^{l}$ is the ground-truth LDR image. The tone-mapped HDR images $\hat{\mathbf{I}}_{t,2D}^{h}$ and $\hat{\mathbf{I}}_{t}^{h}$ are derived from rendered HDR image $\mathbf{I}_{t,2D}^{h}$ and ground-truth HDR image $\mathbf{I}_{t}^{h}$, respectively, using $\mu$-law compression: $\mathbf{\hat{I}}^{h}=\frac{\log(1+\mu\cdot\text{norm}(\mathbf{I}^{h}))}{\log(1+\mu)}$ where $\mu$ is a compression factor and norm()($\cdot$) is min-max normalization. This transformation aligns HDR and LDR domains for consistent comparison. Notably, we choose the LDR images $\mathbf{I}_{t,3D}^{l}$ rasterized by Gaussian Splatting as our final LDR rendering results.
@@ 4 Experiments
Implementation details. HDR-4DGS is trained with the Adam optimizer using the same parameters as 4DGS (Yang et al., 2024). For our dynamic tone mapper, the learning rate is set to $5\times 10^{-4}$, and the dimension of temporal radiance context features is set to 2. We adopt the same structure of per-channel tone-mapping functions in our dynamic tone mapper as HDR-GS (Cai et al., 2024) and GRU (Cho et al., 2014) is adopted to implement the dynamic radiance context learner by default. For equation 7, $\lambda$ is set to 0.2, and $\alpha$ is set to 0.6 if HDR ground truth is available.
Evaluation metrics. We adopt the PSNR and SSIM as quantitative metrics, LPIPS as an additional perceptual metric. Following prior works (Cai et al., 2024; Huang et al., 2022; Liu et al., 2025a), we apply Photomatix Pro (HDRsoft Team, 2025) to convert HDR images into displayable LDR images for qualitative visualization and fair comparison. Futhermore, training time and inference speed (fps) are reported. Results are averaged over all scenes.